[figure: 9.7 cm]

MAGNETISM AND ELECTRONIC STRUCTURE OF FERROMAGNETIC Gd BISMUTHIDES AND SILICIDES

J. Szade, M. Drzyzga and G. Skorek

Institute of Physics, University of Silesia, Uniwersytecka 4, 40-007 Katowice, Poland

Gd_{4}Bi_{3} and Gd_{5}Si_{4} exhibit the highest ferromagnetic order transition (∼335 K) among the Gd compounds with nonmagnetic elements. Their magnetic susceptibility was investigated in a large temperature range and the limited applicability of the Curie-Weiss law was found. A temperature dependent susceptibility of the Gd 5d electrons is suggested as a possible source of the nonlinear inverse susceptibility in the temperature range above T_{C}. We investigated also the Gd_{5}Bi_{3}, which shows 2 magnetic transitions below 270 K. The electronic structure of both bismuthides was studied with use of x-ray photoelectron spectroscopy (XPS). The valence band of Gd_{5}Bi_{3} exhibits, in the vicinity of the Fermi level, similar features to the described earlier antiferromagnetic GdBi, whereas for the ferromagnetic Gd_{4}Bi_{3} the shape of the photoelectron energy distribution curve is different. It shows a maximum at binding energy of about 2 eV and a feature at 0.5 eV. A similar shape was obtained for Gd_{5}Si_{4} and one can relate the ferromagnetism in those compounds with a particular energy distribution of the Gd 5d electrons mediating the coupling between the 4f moments. The chemical shift of the core levels and possible charge transfer were also analysed.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : J. Szade

Address for correspondence : Institute of Physics, University of Silesia Uniwersytecka 4 40-007 Katowice

Email address : szade@us.edu.pl